Sunland Company is constructing a building. Construction began on February 1 and was completed on December 31. Expenditures were $900,000 on March 1, $600,000 on June 1, and $1,500,000 on December 31.

Sunland Company borrowed $500,000 on March 1 on a 5-year, 12% note to help finance construction of the building. In addition, the company had outstanding all year a 12%, 5-year, $1,000,000 note payable and an 11%, 4-year, $1,750,000 note payable. Compute avoidable interest for Sunland Company. Use the weighted-average interest rate for interest capitalization purposes. (Round "Weighted-average interest rate" to 4 decimal places, e.g. 0.2152 and final answer to 0 decimal places, e.g. 5,275.
Answer #1

If it is helpful, please rate the answer and if any doubt arises let me know

Avoidable Interest $                 1,28,182
Workings:
Expenditure for the year
Mar-01 $                 9,00,000 X 10 / 12 = $   7,50,000
Jun-01 $                 6,00,000 X 7 / 12 = $   3,50,000
Dec-31 $              15,00,000 X 0 / 12 = $                -  
$              30,00,000 $ 11,00,000
Interest Capitalized
$                         11,00,000
Less: $                            5,00,000 X 12.00% = $       60,000
Balance $                            6,00,000 X 11.3636% = $       68,182
Interest Capitalized = $   1,28,182
Weighted Average rate of all debt:-
$              10,00,000 X 12% = $   1,20,000
$              17,50,000 X 11% = $   1,92,500
$              27,50,000 $   3,12,500
Weighted Average rate of all debt = 11.3636%
($312500 / $2750000)